# Swamplandish Unification of the Dark Sector

### v1

@@
We provide a short overview of recent progress made in our understanding of the dark sector based on the Swampland program which in turn is rooted in lessons from string theory. We explain how the existence of one extra mesoscopic dimension (the “dark dimension”) in the micron range emerges and how this can lead to a unification of the dark energy and dark matter. In particular the smallness of the dark energy leads to the prediction of the existence of a tower of weakly interacting light particles which can naturally play the role of dark matter. Moreover this unifies dark matter with gravity as dark matter ends up being excitations of graviton in the dark dimension. We also explain how in combination with other Swampland principles one finds an explanation of the “why now” and the “cosmological coincidence” problems. This model is consistent with the cosmological bounds as well as the Newton’s inverse square law, but makes predictions which differ from $\Lambda$CDM. This review is intended for a broad audience of high energy theorists and cosmologists without prior knowledge of string theory and it explains the motivations and predictions of this program in a non-technical form.
@@ 1 Introduction
The main aim of this short note is to explain the relevant parts of the Swampland program which have found surprisingly powerful applications in the context of the dark sector of our universe and in particular summarize the recent work [4, 5, 6, 7] leading to the prediction of the dark dimension and the unification of the dark sector. We also explain some consequences of the de Sitter conjectures [8, 9] for the dark sector. The organization of this paper is as follows: In section 2 we review aspects of the Swampland program relevant for the dark sector. In section 3, we show how applying these principles to our universe leads to the prediction of one extra mesoscopic dimension in the micron range. In section 4 we discuss the cosmological aspects of this model and explain how dark matter emerges in a natural way in this setup. In section 5 we discuss experimental prospects for testing these ideas. We end in section 6 with some concluding thoughts.
@@ 2 Relevant Swampland Principles
Figure 1: As we go to extreme values of field space, a tower of light particles emerge whose mass scales exponentially with field space distance.
@@ 4 Cosmological aspects of the dark dimension scenario
Figure 2: The standard model brane, which is localized in the micron scale 5-th dimension, radiates gravitons into bulk, which serve as KK dark matter. Most of the initial dark matter distribution is at a mass close to the initial temperature.
@@ 5 Experimental constraints
Figure 3: The KK gravitons are not stable and after they are created at a mass mostly around $\sim\rm{GeV}$ decay to lighter KK gravitons, today reaching a mass in the 50-100 $\rm{keV}$ range.
@@ 6 Concluding thoughts
While based on all we have seen it is natural to expect that in the near future we will find exciting evidence for the dark dimension scenario, it is natural to ask what would it mean, theoretically, if the experiments looking for deviations for ISL fail to observe it in the micron range? Here we assess the certainty we have of the validity of various assumptions that go into predicting the dark dimension. In the application of the Swampland’s distance conjecture to dS it was argued based on naturalness that $m\lesssim\Lambda^{1/4}$ [4]. If the exponent in this relation was different then the Swampland principle could still be valid but with a different exponent. However, note that if we believe that the dark matter and dark energy are unified, as it is rather natural in the context of distance conjecture, using the fact that the mass of the dark matter today cannot be smaller than about $100\ \rm{keV}$ (due to extra-galactic background radiation and distortions on CMB) leads to the prediction that $m_{KK}\leq 100\ {\rm keV}$, which implies $m\lesssim\Lambda^{0.2}$, and that the ISL experiments would need to find a deviation at scale no smaller than $10^{-4}\mu m$. If indeed no deviation of ISL is found even at that scale, it could be that the dark matter and dark energy unification does not arise as expected and the exponent of the distant conjecture is even smaller than $0.2$. Another option is that despite the massive amount of evidence for the distance conjecture for $\Lambda<0$, and the theoretical argument for it, which does not depend on the sign of $\Lambda$ perhaps there is a subtlety and it does not extend to $\Lambda>0$. We can only wait for the experimental verdict. Either way, we will learn exciting new physics!

### v2

@@
We provide a short overview of recent progress made in our understanding of the dark sector based on the Swampland program which in turn is rooted in lessons from string theory. We explain how the existence of one extra mesoscopic dimension (the “dark dimension”) in the micron range emerges and how this can lead to a unification of the dark energy and dark matter. In particular the smallness of the dark energy leads to the prediction of the existence of a tower of weakly interacting light particles which can naturally play the role of dark matter. Moreover this unifies dark matter with gravity as dark matter ends up being excitations of graviton in the dark dimension. We also explain how in combination with other Swampland principles one finds an explanation of the “why now” and the “cosmological coincidence” problems. This model is consistent with the cosmological bounds as well as the Newton’s inverse square law, but makes predictions which differ from $\Lambda$CDM. It also gives rise to an appealing picture of hierarchy of scales in particle physics pegged to the dark energy, including a possible origin of the electroweak hierarchy and the prediction of masses of QCD axion and sterile neutrinos both in the 1-10 meV range. This review is intended for a broad audience of high energy theorists and cosmologists without prior knowledge of string theory and it explains the motivations and predictions of this program in a non-technical form.
@@ 1 Introduction
The main aim of this short note is to explain the relevant parts of the Swampland program which have found surprisingly powerful applications in the context of the dark sector of our universe and in particular summarize the recent work [4, 5, 6, 7] leading to the prediction of the dark dimension and the unification of the dark sector. We also explain some consequences of the de Sitter conjectures [8, 9] for the dark sector. We also discuss predictions of this scenario for particle phenomenology, including an appealing picture of scale hierachy. The organization of this paper is as follows: In section 2 we review aspects of the Swampland program relevant for the dark sector. In section 3, we show how applying these principles to our universe leads to the prediction of one extra mesoscopic dimension in the micron range. In section 4 we discuss the cosmological aspects of this model and explain how dark matter emerges in a natural way in this setup. In section 5 we discuss experimental prospects for testing these ideas. We conclude in section 6 with some observations about predictions of this scenario for particle phenomenology.
@@ 2 Relevant Swampland Principles
Figure -359: As we go to extreme values of field space, a tower of light particles emerge whose mass scales exponentially with field space distance.
@@ 4 Cosmological aspects of the dark dimension scenario
Figure -358: The standard model brane, which is localized in the micron scale 5-th dimension, radiates gravitons into bulk, which serve as KK dark matter. Most of the initial dark matter distribution is at a mass close to the initial temperature.
@@ 5 Experimental constraints
Figure -357: The KK gravitons are not stable and after they are created at a mass mostly around $\sim\rm{GeV}$ decay to lighter KK gravitons, today reaching a mass in the 50-100 $\rm{keV}$ range.
@@ 6 Concluding thoughts
This picture also leads to an interesting picture for QCD axion [27]: The axion naturally is localized on the standard model brane which leads to the prediction that the QCD axion decay constant $f_{a}\lesssim\hat{M_{p}}\sim 10^{10}\ GeV$. Combined with observational lower bounds for QCD axion this narrows the axion mass to a small window in the range $1-10\ meV$. Also $\hat{M_{p}}\sim\Lambda^{\frac{1}{12}}\sim 10^{10}\ GeV$ may signal a scale where new physics takes over. For example it could be an explanation of why the Higgs potential has an instability in this scale.
It is intriguing that the neutrino, sterile neutrino, dark tower scale separation and the QCD axion all end up in the same range of about $\Lambda^{\frac{1}{4}}$. Indeed in the dark dimension scenario the various scales in physics are all roughly organized in integral powers of the 5d Planck scale $\Lambda^{\frac{1}{12}}$:
$\qquad\Lambda\sim 10^{-120}M_{p}^{4}$
${\rm H}_{0}\sim\tau_{\text{now}}^{-1}$ $\sim\Lambda^{\frac{6}{12}}\sim 10^{-60}M_{p}\sim 10^{-40}\,\text{GeV}$
$m_{\nu}\sim m_{a}$ $\sim\Lambda^{\frac{3}{12}}\sim 10^{-30}M_{p}\sim 10^{-10}\,\text{GeV}$
$\Lambda_{\text{QCD}}\sim\alpha\Lambda_{\text{weak}}$ $\sim\Lambda^{\frac{2}{12}}\sim 10^{-20}M_{p}\sim\,1\,\text{GeV}$
$\Lambda_{\text{Higgs inst.}}\sim{\hat{M}}_{p}\sim f_{a}$ $\sim\Lambda^{\frac{1}{12}}\sim 10^{-10}M_{p}\sim 10^{10}\,\text{GeV}$ (6.3)
While based on all we have seen it is natural to expect that in the near future we will find exciting evidence for the dark dimension scenario, it is important to ask what would it mean, theoretically, if the experiments looking for deviations for ISL fail to observe it in the micron range? Here we assess the certainty we have of the validity of various assumptions that go into predicting the dark dimension. In the application of the Swampland’s distance conjecture to dS it was argued based on naturalness that $m\lesssim\Lambda^{1/4}$ [4]. If the exponent in this relation was different then the Swampland principle could still be valid but with a different exponent. However, note that if we believe that the dark matter and dark energy are unified, as it is rather natural in the context of distance conjecture, using the fact that the mass of the dark matter today cannot be smaller than about $100\ \rm{keV}$ (due to extra-galactic background radiation and distortions on CMB) leads to the prediction that $m_{KK}\leq 100\ {\rm keV}$, which implies $m\lesssim\Lambda^{0.2}$, and that the ISL experiments would need to find a deviation at scale no smaller than $10^{-4}\mu m$. If indeed no deviation of ISL is found even at that scale, it could be that the dark matter and dark energy unification does not arise as expected and the exponent of the distant conjecture is even smaller than $0.2$. Another option is that despite the massive amount of evidence for the distance conjecture for $\Lambda<0$, and the theoretical argument for it, which does not depend on the sign of $\Lambda$ perhaps there is a subtlety and it does not extend to $\Lambda>0$. We can only wait for the experimental verdict. Either way, we will learn exciting new physics!